OpenStudy (anonymous):

Newton’s universal law of gravitation says that the force of gravitational attraction, F between two particles is directly proportional to the product of their masses, 1 m and 2 m , and inversely proportional to the square of the distance, d , between them. If the two particles each lose half of their mass as the distance between them triples, how would the gravitational attraction between them change

OpenStudy (anonymous):

\[-G \frac{Mm}{r^2}\]

OpenStudy (anonymous):

so let's do what problem say \[-G \frac{(1/2M)(1/2m)}{(3r)^2}\]

OpenStudy (anonymous):

\[-G \frac{(1/2M)(1/2m)}{(3r)^2}=-G \frac{1/4(M)(m)}{9(r)^2}=-G ((1/4) /9)\frac{(M)(m)}{(r)^2}\]

OpenStudy (anonymous):

can you wxplain please

OpenStudy (anonymous):

I took the gravity equations \[-G \frac{Mm}{r^2}\] and your problem states "If the two particles each lose half of their mass as the distance between them triples," so that's what I did \[-G \frac{(1/2M)(1/2m)}{(3r)^2}\] and I simplified it \[-G ((1/4) /9)\frac{(M)(m)}{(r)^2}= \frac{1}{36}(-G \frac{Mm}{r^2} )\]

OpenStudy (anonymous):

so gravitation force change by 1/36

OpenStudy (anonymous):

Better now sorry its just confusing. great work

OpenStudy (anonymous):

so main thing is changing each part and noticing how it affects the whole thing